On January 1, 2018, the Shagri Company began construction on a new manufacturing facility for its own use. The building was completed in 2019. The only interest-bearing debt the company had outstanding during 2018 was long-term bonds with a book value of $11,700,000 and an effective interest rate of 8%. Construction expenditures incurred during 2018 were as follows:

January 1 $ 670,000
March 1 702,000
July 31 582,000
September 30 770,000
December 31 470,000


Required:
Calculate the amount of interest capitalized for 2018.

Date Expenditure Weight Average
January 1 x =
March 1 x =
July 31 x =
September 30 x =
December 31 x =
Accumulated expenditure $0 $0
Average Interest Rate Capitalized Interest
Average accumulated expenditures $0 x % = $0
Answer #1

Calculate the amount of interest capitalized for 2018.

Date Expenditure * Weight = Average
January 1 670000 * 12/12 = 670000
March 1 702000 * 10/12 = 585000
July 31 582000 * 5/12 = 242500
September 30 770000 * 3/12 = 192500
December 31 470000 * 0/12 = 0
Accumulated expenditure 3194000 1690000
Average * Interest rate = Capitalized interest
Average accumulated expenditure 1690000 * 8% = 135200